Q 13.32: A wall is built of two metals (stainless steel and lead) that are in thermal contact with each other. If the temperature difference across both metals is 28.856 K and across the (3.001 cm thick) lead plate is 12.32 K, how thick is the steel plate?
Answer #1

Ab teel lolead. 3 Noming on alls is done to eabily write temperature ex poebbion. Ti – 13 = 28.856K -O Te - T3 = 12.32 k to †1. So, I steel = t totalt lead = (7.0289 - 3.001) cm. to tell = 4.02 79 cm Anse, So, thickness of steel plate P8 4.0279 cm